16g of SOx occupies 5.6 litre volume at S. T. P assuming ideal gas nature the value of X is

Grade:10

3 Answers

Utkarsh Mishra
6 years ago
At STP one mole of a gas occupies 22.4 l of volumeso no of moles=given volume/22.4 =5.6/22.4 =1/4 molesA/Q16g of SOx =1/4 moleso,weight of one mole =64 gor,Moor mass of SOx =64gso,32g+16x g=64ghence, x=2
Anjali kumari
6 years ago
As we know volume at STP=22.4LTherefore no. of moles=5.6L/22.4L=1/4 moles.Now, no. of moles= given mass/ molecular mass. Hence,1/4=16/molecular mass. Or,molecular mass=64g.This means that molecular mass of SOx=64g.Now molecular mass of SOx=32+X=64.ThereforeX=32.So, the value of x=2.
Rohit
6 years ago
First we need to get the no. of moles of the gas. It is n= V/22.4...... u will get it as 0.25 moles...now n=(given weight)/molar mass...........hence molar mass=1600/25..... The molar mass is 64g ....... Now the at. Wt of S is 32.....hence 32+x=64. At wt of O is 16 hence we need 2 O`s in order to get the given molar mass.Thanks Regards ROHIT